OpenStudy (anonymous):

Water coming out from a fountain is modeled by the function f(x) = -x2 + 5x + 4 where f(x) represents the height, in feet, of the water from the fountain at different times x, in seconds. What does the average rate of change of f(x) from x = 3 to x = 5 represent?

ganeshie8 (ganeshie8):

rate of change of f(x) between x = a and x = b : \[\large \dfrac{f(b)-f(a)}{b-a}\]

ganeshie8 (ganeshie8):

a = 3 b = 5 evaluate f(3), f(5) and plugin

OpenStudy (anonymous):

I'm confused @ganeshie8

OpenStudy (anonymous):

So (5f - 3f)/5-3 ?

ganeshie8 (ganeshie8):

nice try, but no. you need to evaluate : \[\large \dfrac{f(5) - f(3)}{5-3}\]

ganeshie8 (ganeshie8):

f(5) is NOT same as 5f f(5) is a notation which represents the function f(x) evaluated at x = 5

ganeshie8 (ganeshie8):

you have : f(x) = -x^2 + 5x + 4 can you evaluate f(5) ? f(5) = ?

OpenStudy (anonymous):

Ohhh I see now

OpenStudy (anonymous):

Give me an minute

OpenStudy (anonymous):

f(5) = 54

OpenStudy (anonymous):

f(3) = 28

OpenStudy (anonymous):

54 -28 / 5 - 3 = 13

ganeshie8 (ganeshie8):

careful, its -x^2

OpenStudy (anonymous):

But negative times a negative is positive right?

OpenStudy (anonymous):

So it wouldn't matter if the negative is there

ganeshie8 (ganeshie8):

f(x) = -x^2 + 5x + 4 f(5) = -(5^2) + 5(5) + 4 = -25 + 25 + 4 = 0 + 4 = 4

ganeshie8 (ganeshie8):

f(3) = -(3^2) + 5(3) + 4 = -9 + 15 + 4 = 10

ganeshie8 (ganeshie8):

the average rate would be : \[\large \dfrac{4 - 10}{5-3}\] simplify

OpenStudy (anonymous):

-3

ganeshie8 (ganeshie8):

\(\large \color{red}{\checkmark}\)

ganeshie8 (ganeshie8):

Notice that by P `E`MDAS, Exponent should be evaluated before Multiplication, when you have an expression like `-x^2`

ganeshie8 (ganeshie8):

-x^2 is not same as (-x)^2

OpenStudy (anonymous):

So what does the average rate of change of f(x) from x = 3 to x = 5 represent?

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

what do you think it represents ?

ganeshie8 (ganeshie8):

\[\large \text{Average rate of change}=\dfrac{f(b)-f(a)}{b-a}\]

OpenStudy (anonymous):

That the water comes out of the fountain at a speed of 3 feet per second

ganeshie8 (ganeshie8):

numerator has f(x) whose units are `feet` right?

ganeshie8 (ganeshie8):

and the denominator has units of seconds `seconds`

ganeshie8 (ganeshie8):

so the units of average rate of change would be `feet/sec` which is same as the units of speed, so yeah you're right !

OpenStudy (anonymous):

Awesome, thanks